If a plane travels 1440km in 2 hours. How fast is the plane travelling in metres per second.
Answer by stanbon(75887)

If a plane travels 1440km in 2 hours. How fast is the plane travelling in metres per second.
-------
1440/2 (km/hr) = 770 (km/hr)(1000 m/1km)(1 hr/3600 sec)
--------------
= (770*1000*1)/(1*1*3600)(meters/second)
=====
= 213.89 meters/sec
=======================
Cheers,
Stan H.
=======================